*imaginary no. problem*
z = -3-9i
write the complex conjugate no. z and find zz. (sorry, I cannot put the line on top of the second z.)

Answer by oscargut(2103)

z conjugate is: -3+9i
z(z conjugate) = (-3-9i)(-3+9i) = 9+81 =90
I can help you with more questions at: mthman@gmail.com
Thanks